# Online Posterior Sampling with a Diffusion Prior

### v1

@@ C.3 Ablation Studies
Algorithm 4 $\tt DPS$ of Chung et al. [12].
@@
5:     $\hat{S}\leftarrow\frac{\varepsilon_{t}(S_{t};\psi_{t})}{\sqrt{1-\bar{\alpha}_{t}}}$
6:     $\hat{S}_{0}\leftarrow\frac{1}{\sqrt{\bar{\alpha}_{t}}}(S_{t}+(1-\bar{\alpha}_{t})\hat{S})$
@@
In our experiments, we compare to diffusion posterior sampling ($\tt DPS$) with a Gaussian observation noise (Algorithm 1 in Chung et al. [12]). We implement it as shown in Algorithm 4. The score is $\hat{S}=\varepsilon_{t}(S_{t};\psi_{t})/\sqrt{1-\bar{\alpha}_{t}}$ because we regress $\tilde{\varepsilon}_{t}$ in Appendix B. We set $\tilde{\sigma}_{t}=\sqrt{\tilde{\beta}_{t}}$, which is the same noise as in our reverse process (Section 3). The term
@@
is the gradient of the negative log-likelihood with respect to $S_{t}$ evaluated at $\hat{S}_{0}$.
As noted in Appendices C.2 and D.1 of Chung et al. [12], $\zeta_{t}$ in $\tt DPS$ needs to be tuned for good performance. We also observed this in our experiments (Section 6.2). To make $\tt DPS$ work well, we follow Chung et al. [12] and set
@@
While this significantly improves the performance of $\tt DPS$, it does not prevent failures. The fundamental problem is that gradient-based optimization is sensitive to the step size, especially when the optimized function is steep. We do not have any such hyper-parameter in our algorithm.

### v2

@@ C.4 Non-Bandit Evaluation
Algorithm 4 $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ of Chung et al. [12].
@@
5:  $\hat{S}\leftarrow-\frac{\varepsilon_{t}(S_{t};\psi_{t})}{1-\bar{\alpha}_{t}}$
[…]
@@
In our experiments, we compare to diffusion posterior sampling ($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$) with a Gaussian observation noise (Algorithm 1 in Chung et al. [12]). Our implementation is presented in Algorithm˜4. The score is $\hat{S}=-\varepsilon_{t}(S_{t};\psi_{t})/(1-\bar{\alpha}_{t})$, where $\varepsilon_{t}(S_{t};\psi_{t})$ is a regression estimate of the forward process noise $\varepsilon_{t}$ in Appendix˜B. We set $\tilde{\sigma}_{t}=\sqrt{\tilde{\beta}_{t}}$, which is the same amount of noise as in our reverse process (Section˜3). The term
@@
is the gradient of the negative log-likelihood with respect to $S_{t}$.
As discussed in Appendices C.2 and D.1 of Chung et al. [12], $\zeta_{t}$ in $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ needs to be tuned for good performance. This is because $\nabla\sum_{\ell=1}^{N}(y_{\ell}-\phi_{\ell}^{\top}\hat{S}_{0})^{2}$ grows with the number of observations, which causes instability. We also observed this in our experiments (Section˜6.2). To make $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$ work well, we follow Chung et al. [12] and set
@@
While this significantly improves the performance of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DPS$, it does not prevent failures. The fundamental problem is that gradient-based optimization is sensitive to the step size, especially when the optimized function is steep. Note that $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt LaplaceDPS$ does not have any such hyper-parameter.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state all contributions. The introduction also points to where those contributions are made.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: The increase in computational cost is discussed in Section˜4.2 and shown empirically in Section˜6.2. We also conduct an ablation study in Section˜C.3, where we show how the regret of $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$ scales with the number of samples used for pre-training the prior and the number of stages in the diffusion model prior.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [Yes]
  12. Justification: The main claims are stated and discussed in Section˜4. Their proofs are in Appendix˜A.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: We also include code to reproduce the synthetic results in Figures˜2 and 4.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: We include code to reproduce the synthetic results in Figures˜2 and 4.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: The experiments are described to a sufficient level to be reproducible. To make sure, we include code to reproduce the synthetic results in Figures˜2 and 4.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: All plots in the paper have error bars.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [No]
  32. Justification: Our experiments are not large scale.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We checked the link and comply.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [N/A]
  40. Justification: This work is algorithmic and not tied to a particular application that would have immediate negative impact.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: This paper does not pose such a risk.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: All used assets are stated and cited.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [N/A]
  52. Justification: This paper does not release new assets.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: No crowdsourcing or research with human subjects.
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: No crowdsourcing or research with human subjects.